OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 3) and is perpendicular to the line y = –x – 3

OpenStudy (anonymous):

ok you are going to need the point slope formula which is y-y1=m(x-x1) so we know our m value or slope is going to be the recipical of -1 which is 1 y-3=1(x-2) y=x+1

OpenStudy (anonymous):

y-3= 1(0-2)

OpenStudy (aroub):

y=mx+b , m= 1 , (2,3) 3=1(2)+b 3=2+b 1=b y=x+1

OpenStudy (anonymous):

yint=1 y=x+1 (slope is positive because of negative recipricoal)